Suppose, in an economy, currency in circulation (C) is $16 billions, reserves (R) held by banks are $4 billions, and deposits (D) by people and firms in banks are worth $ 84 billions. If there are no excess reserves, then

(a) What is the money supply (M) in the economy? _______________

(b) What is the monetary base (MB)? _______________

(c) What is the currency deposit ratio ? _______________

(d) What is the reserve deposit ratio? _______________

(e) What is the money multiplier (m)? _______________

(f) If the monetary base increases by 10%, then the money supply increases by $ _________ billions.

(g) In general, how can the central bank influence the monetary base?
Answer #1

Answer

(a) Money Supply(M) = C + D = 16 billion + 84 billion= $100 billion

(b) Monetary Base = C + R = 16 billion + 4 billion + 0 = $20 billion

(c) currency deposit ratio(cr) = C/D = (16 billion) / (84 billion) = 4/21 = 0.19(approx)

(d) Reserve deposit rate(rr) = RR/D = (4 billion)/(84 billion) = 1/21 = 0.0476(approx) (Note that there are no eacess reserves, thus R = RR + ER and ER = excess reserves = 0. thus RR = required reserves = R)

(e) Money multiplier = (1 + cr)/(cr + rr + er) where er = excess reserve ratio = 0 as excess reserves are 0.

=> Money multiplier = (1 + 4/21)/(4/21 + 1/21) = 5

=> Money multiplier = 5

(f) 10% of monetary base = 10% of 20 billion = 2 billion.

Thus new monetary base = 22 billion.

Money multiplier is 5.

Thus New Money supply = money multiplier*monetary base = 22 billion*5 = 110 billion.

Thus money supply will increase by 110 billion - 100 billion = $10 billion

(g) Central bank influences by changing reserves for the banks and they do this by using open market operations like buying and selling bonds. They, can also do this by changing bank rate.